The Bruin's Den Outdoor Gear is considering a new 7-year project to produce a new tent line. The equipment necessary would cost $1.63 million and be depreciated using straight-line depreciation to a book value of zero. At the end of the project, the equipment can be sold for 15 percent of its initial cost. The company believes that it can sell 26,500 tents per year at a price of $70 and variable costs of $30 per tent. The fixed costs will be $455,000 per year. The project will require an initial investment in net working capital of $217,000 that will be recovered at the end of the project. The required rate of return is 11.3 percent and the tax rate is 35 percent. What is the NPV?


           
            A)    $1,203,997         
            B)    $660,711
            C)    $823,854
            D)    $507,099
            E)    $546,275

Please show work, we are not allowed to use excel on our exam
Answer #1

Depreciation = Cost of the Project / Useful life years = $1,630,000 / 7 = $232,857.14

Cash flow at year 0 = Cost of Project + Investment in NWC = $1,630,000 + $217,000 = $1,847,000

Annual Cash Flow(Year 1 - 7) = [{(P - VC) * Qty. Sold} - Fixed Cost] * [1 - t] + [Depreciation * t]

= [{($70 - $30) * 26,500} - $455,000] * [1 - 0.35] + [$232,857.14 * 0.35]

= $393,250 + $81,500 = $474,750

Additional Cash Flow at year 7 = After-tax Salvage Value + Recovery of NWC

= [($1,630,000 * 0.15) * (1 - 0.35)] + $217,000 = $158,925 + $217,000 = $375,925

NPV = PV of Cash Inflows - PV of Cash Outflows

= [$474,750 * {1 - (1 + 0.113)-7} / 0.113] + [$375,925 / (1 + 0.113)7] - $1,847,000

= [$474,750 * 4.6669] + [$375,925 / 2.1158] - $1,847,000

= $2,215,596.60 + $177,678.57 - $1,847,000 = $546,275.17

So, Option "E" is correct.